OpenStudy (anonymous):

Solve for x. 3(x + 5) - 2(x + 2) = 20 1 9 13

OpenStudy (anonymous):

I got 8 the first time and not sure why

OpenStudy (xxdragoonxx):

9.

OpenStudy (anonymous):

Well could you expian how you got that?

OpenStudy (anonymous):

x=20-15+4=9

OpenStudy (uri):

3(x + 5) - 2(x + 2) = 20 3x+15-2x-4=20 3x-2x=20-15+4 x=9

OpenStudy (anonymous):

3x-2x=x

OpenStudy (anonymous):

Thanks, @uri

OpenStudy (anonymous):

Not sure why I got stumped.....

OpenStudy (uri):

Yw,silly mistakes happen. :)

OpenStudy (anonymous):

:) Yes, yes they do! :)